Project

General

Profile

Joint MICE Experiment Management Office and MICE International Project Office

18th September 2018; 15:30 BST

Phone details

Agenda

  1. Introduction: KL
  2. Notes of previous meeting and status of actions: All
    • KL: will migrate EB and CB mailing lists to jiscmail
    • CR: will migrate analysis mailing list to jiscmail
    • PF: will migrate software mailing list to jiscmail
  3. Project Manager's report: CW
  4. Software and computing status: PF
  5. Paper status and plans: CR
  6. CM52: status of planning etc.: KL/All
  7. AOB

Notes

Present: P. Soler, C. Whyte, C. Rogers, K. Long, A. Bross,

  1. Introduction: KL
  2. Notes of previous meeting and status of actions: All
    • KL: will migrate EB and CB mailing lists to jiscmail
      • Stands
    • CR: will migrate analysis mailing list to jiscmail
      • Stands
    • PF: will migrate software mailing list to jiscmail
      • Stands
  3. Project Manager's report: CW
    • All detectors are now shipped. Mississippi colleagues have removed the aerogel.
    • MICE Hall decommissioning continues: H2-gas panel and hydride bed has been removed. South Mezzanine mods are underway.
    • PPD ownership now for experiment. Hole in synchrotron wall will be blocked on the inside so that there are no more any RP issues.
    • Quads are being removed, most likely they will be scrapped unless a home can be found. These quads need de-ionized water and have some earth-leakage issues. PS will contact CERN East Area Test Beams to see if there is interest.
    • Focus coils will be stored, being organised by J. Boehm.
    • PS brought up issie of recovery of TDC purchased for MICE. CW will organise recovery of boards for CM52.
    • Organise a walk-around in the MICE Hall to see whether there is any equipment still worth recovering (action CW).
    • Tracker Lab is to be repurposed; need to plan recovery of the equipment. (action KL/ABr)
  4. Software and computing status: PF
    • No Paolo.
  5. Paper status and plans: CR
    • Tracker analysis/reconstruction: Considering a noise-reduction algorithm based on FADC count or NPE. This will then be used to inform improvements in pattern recognition.
    • Emittance-measurement paper: Revisit of consideration of apertures did not yield an apperture. Now moving 'experiment' to correct analysis. Indications are that this will help.
    • Emittance-evolution paper: Referees meeting tomorrow. Awaiting result of this meeting.
    • System-performance paper: No update since meeting on paper last week. CM52 we are expecting a draft of the paper.
    • 6D/wedge paper: C. Brown, following LH2 analysis, will pick up this analysis.
    • Scattering: J.Nugent plans to update this week. PS reported that following resolution of computing issues, JN is re-evaluating the alignment.
    • LH2 scattering: G. Chapzitheodoridis is now progressing the LH2 scattering analysis. LiH results have been reproduced and he intends to present his results at the CM.
    • Field-on scattering: A. Young continues to work on this.
    • Students have been at RAL Summer School, so limitted progress in recent past.
    • LH2 volume uniformity of density has been discussed with J. Boehm and C. MacWaters. Mis-calibration may have been introduced by current limiters used to limit degree of heating produced by level sensors in LH2.
  6. CM52: status of planning etc.: KL/All
    • KL will send email reminding people that the accommodation block booking is almost saturated and will be dropped tomorrow.
  7. AOB
    • Pressure on office space is mounting. KL will email collaboraiton to say need to clear so that we can reduce space use by factor of 2 on timescale of Christmas.
    • D. Rajaram back for three months. Working on track reconstruction with a view to making sure we dont have to look at it again when he is done. We agree to invite him to MEMO/EB.

Summary of actions:

  • KL: will migrate EB and CB mailing lists to jiscmail
  • CR: will migrate analysis mailing list to jiscmail
  • PF: will migrate software mailing list to jiscmail
  • CW: organise recovery of TDC boards at CM52.
  • CW: set up 'walk around' of Hall at CM52 to identify remaining valuable equipment.
  • KL/ABr: Organise recovery of tracker equipment from tracker lab at CM52.
  • KL: Remind collaboration of imperative to book accommodation for CM52 asap.
  • KL: Inform collaboraiton of impending reduction in office space.
  • KL: Invite DR to MEMO/EB.

Back to MEMO:

Back

Back to MIPO:

Back

Documents